# RLP: Reinforcement as a Pretraining Objective

### v1

@@ Abstract
The dominant paradigm for training large reasoning models starts with pre-training using next-token prediction loss on vast amounts of data. Reinforcement learning, while powerful in scaling reasoning, is introduced only as the very last phase of post-training, preceded by supervised fine-tuning. While dominant, is this an optimal way of training? In this paper, we present RLP, an information-driven reinforcement pretraining objective, that brings the core spirit of reinforcement learning—exploration—to the last phase of pretraining. The key idea is to treat chain-of-thought as an exploratory action, with rewards computed based on the information gain it provides for predicting future tokens. This training objective essentially encourages the model to think for itself before predicting what comes next, thus teaching an independent thinking behavior earlier in the pretraining. More concretely, the reward signal measures the increase in log-likelihood of the next token when conditioning on both context and a sampled reasoning chain, compared to conditioning on context alone. This approach yields a verifier-free dense reward signal, allowing for efficient training for the full document stream during pretraining. Specifically, RLP reframes reinforcement learning for reasoning as a pretraining objective on ordinary text, bridging the gap between next-token prediction and the emergence of useful chain-of-thought reasoning. Pretraining with RLP on qwen3-1.7b-base lifts the overall average across an eight‑benchmark math‑and‑science suite by 19%. With identical post‑training, the gains compound, with the largest improvements on reasoning‑heavy tasks such as AIME25 and MMLU‑Pro. Applying RLP to the hybrid Nemotron-Nano-12B-v2 increases the overall average from 42.81% to 61.32% and raises the average on scientific reasoning by 23%, demonstrating scalability across architectures and model sizes. Code: https://github.com/NVlabs/RLP
@@ 2 Methodology
We introduce RLP, a pretraining-time procedure that explicitly induces reasoning. As illustrated in Fig.˜1, RLP inserts a short Chain-of-Thought (CoT) before next-token prediction and measures how much that thought improves the model’s log-probability of the observed token relative to a no-think baseline. This improvement, which is a log-likelihood ratio, is a verifier-free, dense reward available at every position in ordinary text corpora. By valuing thoughts in proportion to their predictive benefit, RLP turns reinforcement pretraining into learning to think on the same data used for standard next-token training.
@@ 3 Experimental Setup
We experiment with qwen3-1.7b-base (Yang et al., 2025) and then scale our experiments to a larger Nemotron-Nano-12B-v2 (Nano, 2025) model. [^1] Details about hyper-parameters for each of the below phases can be found in §9.
@@
We apply RLP on a diverse set of datasets across two settings: (i) SFT-style reasoning corpora, including a math-centric set (OmniMath (Gao et al., 2024)) and mixed math + general-reasoning sets (OpenThoughts (Guha et al., 2025), Nemotron-Crossthink (Akter et al., 2025)); and (ii) general-purpose pretraining corpora, covering academic papers (ACAD), math textbooks (Math-Text), and open-ended web pages QA pairs from Common Crawl (Web-Crawl)(Nano, 2025). We train with RLP for 1B tokens using general pretraining corpora ($\mathcal{D}_{\mathrm{PT}}$) to evaluate its effect in an end-to-end llm pretraining pipeline. We denote this model as $\mathcal{M}_{\textsc{RLP}}$.
@@ 4 Results
Table 1 reports the performance of qwen3-1.7b-base under different pretraining and post-training objectives. First, RLP consistently outperforms both the $\mathcal{M}_{\mathrm{base}}$ and $\mathcal{M}_{\mathrm{CPT}}$ across nearly all benchmarks, with especially strong gains on reasoning-heavy tasks such as AIME25 and MMLU-Pro. We see that $\mathcal{M}_{\textsc{RLP}}$ is relatively on average 19% and 17% better than $\mathcal{M}_{\mathrm{base}}$ and $\mathcal{M}_{\mathrm{CPT}}$ respectively. This highlights the effectiveness of dense, verifier-free reinforcement signals for instilling reasoning capabilities during pretraining.
Model Math Avg Science Avg Science Avg@1[4] Avg
$\mathcal{M}_{\mathrm{base}}$ 61.38 34.51 32.54 42.81
$\mathcal{M}_{\textsc{RLP}}$ 65.33 57.26 61.37 61.32
Table 2: Comparison of $\mathcal{M}_{\mathrm{base}}$ and RLP for the Nemotron-Nano-12B-v2 model.
@@
Second, the benefits of RLP persist even after strong post-training (SFT + RLVR). While all models improve after post-training, $\mathcal{M}_{\textsc{RLP}}$ achieves the highest scores with the overall average substantially higher than both $\mathcal{M}_{\mathrm{base}}$ by 8% and $\mathcal{M}_{\mathrm{CPT}}$ by 7% relatively. This indicates that RLP establishes robust reasoning foundations that are not washed out by downstream alignment but instead compound with post-training. We observe particularly large gains in science domains, with $\mathcal{M}_{\textsc{RLP}}$ +Post achieving +3+3 points over $\mathcal{M}_{\mathrm{CPT}}$ +Post. This trend suggests that RLP is not limited to mathematical reasoning but also generalizes effectively to other domains. The ability to strengthen performance in science benchmarks highlights that RLP fosters a broader class of multi-step explanation-driven reasoning skills, moving beyond domain-specific improvements and pointing toward a more versatile foundation for reasoning in LLMs. Overall, the results demonstrate that RLP not only induces reasoning ability during pretraining but also synergizes with post-training, leading to models with stronger and more durable reasoning abilities than those trained with next-token prediction or continuous pretraining.
Scaling Model Size and Architecture:
We further scale RLP to Nemotron-Nano-12B-v2 (Nano, 2025) ($\mathcal{M}_{\mathrm{base}}$), a hybrid Mamba-Transformer language model of 12B parameter size. In this comparison we take an intermediate checkpoint of Nemotron-Nano-12B-v2 trained till $19.8$ trillion tokens and apply RLP for $250$ million tokens only. $\mathcal{M}_{\mathrm{base}}$ on the other hand is trained for $20$ trillion tokens. Table 2 demonstrates that the benefits of RLP persist and even amplify when scaling to larger model sizes and generalizes to different model architectures. On the Nemotron-Nano-12B-v2 model, $\mathcal{M}_{\textsc{RLP}}$ substantially outperforms $\mathcal{M}_{\mathrm{base}}$ across all domains, and particularly $\mathcal{M}_{\textsc{RLP}}$ is relatively 3535% on average better than $\mathcal{M}_{\mathrm{base}}$ inspite of being trained on approx. $200$ billion less tokens. While math performance improves moderately, the most striking gains emerge in science reasoning, where Science Avg jumps an absolute 23%. These results highlight that RLP yields not only stronger math performance but also robust cross-domain reasoning capabilities.
@@
Following the experimental setup in RPT with Omni‑MATH dataset, we
@@
Table 3: RLP outperforms RPT across all averages. qwen3-1.7b-base was trained with both RPT and RLP for one epoch with matched data and compute.
trained both methods for one epoch under matched data and compute budgets before evaluating on our benchmark suite. As summarized in Table Table 3, RLP achieves uniformly higher aggregates: Math Avg improves from 47.50 to 49.62 (+2.12; +4.5% relative), Science Avg from 35.88 to 37.07 (+1.19; +3.3%), and Overall Avg from 41.69 to 43.35 (+1.66; +4.0%). Methodologically, RPT applies reinforcement only to tokens pre‑selected by an auxiliary assistant via entropy filtering and optimizes a sparse, binary next‑token correctness signal that ignores the CoT content, limiting where the signal can be applied. In contrast, RLP evaluates each sampled CoT by the information gain it provides for the observed next token and updates at all positions without an auxiliary filter which yields consistently better averages under the matched setting above. Crucially, this dense, per‑token information‑gain reward supplies richer credit assignment than RPT’s sparse binary signal and, in our matched experiments, empirically yields better performance.
@@ 5 Ablations
Fig.˜2 visualizes the trends across three settings: (a) rollouts, (b) completion length, and (c) KL. Please look into §10 for more detailed numbers and per-task breakdowns. More rollouts help up to $G=16$ (Overall $42.17\%$); $G=4$ and $8$ already reach $41.38\%$ and $41.95\%$, while $G=32$ decreases slightly to $41.75\%$ (Fig. 2a). Increasing completion length gives the largest gains. Specifically, overall rises from $11.50\%$ at $64$ to $42.17\%$ at $2048$, with Math/Science moving from $1.12\%/21.88\%$ to $48.06\%/36.29\%$ (Fig. 2b). Extending to $4096$ yields $42.21\%$ at roughly twice the thought budget, so we default to $2048$. Furthermore, a KL anchor does not help. Specifically, $\beta=10^{-4}$ and $10^{-3}$ give $41.35\%$ and $41.44\%$, compared to $42.17\%$ at $\beta=0$, and it also increases memory and step time (Fig. 2c). We therefore use $G=16$, completion length $2048$, and $\beta=0$ in later experiments.

### v2

@@ Abstract
The dominant paradigm for training large reasoning models starts with pre-training using next-token prediction loss on vast amounts of data. Reinforcement learning, while powerful in scaling reasoning, is introduced only as the very last phase of post-training, preceded by supervised fine-tuning. While dominant, is this an optimal way of training? In this paper, we present RLP, an information-driven reinforcement pretraining objective, that brings the core spirit of reinforcement learning—exploration—to the last phase of pretraining. The key idea is to treat chain-of-thought as an exploratory action, with rewards computed based on the information gain it provides for predicting future tokens. This training objective essentially encourages the model to think for itself before predicting what comes next, thus teaching an independent thinking behavior earlier in the pretraining. More concretely, the reward signal measures the increase in log-likelihood of the next token when conditioning on both context and a sampled reasoning chain, compared to conditioning on context alone. This approach yields a verifier-free dense reward signal, allowing for efficient training for the full document stream during pretraining. Specifically, RLP reframes reinforcement learning for reasoning as a pretraining objective on ordinary text, bridging the gap between next-token prediction and the emergence of useful chain-of-thought reasoning. Pretraining with RLP on qwen3-1.7b-base lifts the overall average across an eight‑benchmark math‑and‑science suite by 19%. With identical post‑training, the gains compound, with the largest improvements on reasoning‑heavy tasks such as AIME25 and MMLU‑Pro. Applying RLP to the hybrid Nemotron-Nano-12B-v2 increases the overall average from 42.81% to 61.32% and raises the average on scientific reasoning by 23%, demonstrating scalability across architectures and model sizes.
@@ 2 Methodology
We introduce RLP, a pretraining-time procedure that explicitly induces reasoning. As illustrated in Fig. 1, RLP inserts a short Chain-of-Thought (CoT) before next-token prediction and measures how much that thought improves the model’s log-probability of the observed token relative to a no-think baseline. This improvement, which is a log-likelihood ratio, is a verifier-free, dense reward available at every position in ordinary text corpora. By valuing thoughts in proportion to their predictive benefit, RLP turns reinforcement pretraining into learning to think on the same data used for standard next-token training.
@@ 3 Experimental Setup
We experiment with qwen3-1.7b-base (Yang et al., 2025) and then scale our experiments to a larger Nemotron-Nano-12B-v2 (Nano, 2025) model. [^1] Details about hyper-parameters for each of the below phases and the prompt used for RLP can be found in Appendix 10.
@@
We apply RLP on a diverse set of datasets across two settings: (i) SFT-style reasoning corpora, including a math-centric set (OmniMath (Gao et al., 2024)) and mixed math + general-reasoning sets (OpenThoughts (Guha et al., 2025), Nemotron-Crossthink (Akter et al., 2025)); and (ii) general-purpose pretraining corpora, covering academic papers (ACAD), math textbooks (Math-Text), and open-ended web pages QA pairs from Common Crawl (Web-Crawl)(Nano, 2025). We train with RLP for 1B input tokens using general pretraining corpora ($\mathcal{D}_{\mathrm{PT}}$) to evaluate its effect in an end-to-end llm pretraining pipeline. We denote this model as $\mathcal{M}_{\textsc{RLP}}$. Note that theoretically RLP can be applied for every token in a document but in our experiments we randomly select one token per document. Hence, the number of tokens for which the reward is applied is far less than 1B.
@@ 4 Results
Table 1 reports the performance of qwen3-1.7b-base under different pretraining and post-training objectives. First, RLP consistently outperforms both the $\mathcal{M}_{\mathrm{base}}$ and $\mathcal{M}_{\mathrm{CPT}}$ across nearly all benchmarks, with especially strong gains on reasoning-heavy tasks such as AIME25 and MMLU-Pro. We see that $\mathcal{M}_{\textsc{RLP}}$ is relatively on average 19% and 17% better than $\mathcal{M}_{\mathrm{base}}$ and $\mathcal{M}_{\mathrm{CPT}}$ respectively. This highlights the effectiveness of dense, verifier-free reinforcement signals for instilling reasoning capabilities during pretraining. Second, the benefits of RLP persist even after strong post-training (SFT + RLVR). While all models improve after post-training, $\mathcal{M}_{\textsc{RLP}}$ achieves the highest scores with the overall average substantially higher than both $\mathcal{M}_{\mathrm{base}}$ by 8% and $\mathcal{M}_{\mathrm{CPT}}$ by 7% relatively. This indicates that RLP establishes robust reasoning foundations that are not washed out by downstream alignment but instead compound with post-training. We observe particularly large gains in science domains, with $\mathcal{M}_{\textsc{RLP}}$ +Post achieving +3+3 points over $\mathcal{M}_{\mathrm{CPT}}$ +Post. This trend suggests that RLP is not limited to mathematical reasoning but also generalizes effectively to other domains. The ability to strengthen performance in science benchmarks highlights that RLP fosters a broader class of multi-step explanation-driven reasoning skills, moving beyond domain-specific improvements and pointing toward a more versatile foundation for reasoning in LLMs. Overall, the results demonstrate that RLP not only induces reasoning ability during pretraining but also synergizes with post-training, leading to models with stronger and more durable reasoning abilities than those trained with next-token prediction or continuous pretraining.
Benchmark $\mathcal{M}_{\mathrm{base}}$ $\mathcal{M}_{\textsc{RLP}}$ $\mathcal{M}_{\mathrm{base}}$ +Post $\mathcal{M}_{\textsc{RLP}}$ +Post MATH500 $79.95$ $78.68$ $83.47$ $87.05$ GSM8K $72.31$ $85.98$ $94.22$ $94.90$ AMC23 $70.63$ $57.19$ $62.19$ $75.00$ Minerva $22.61$ $39.48$ $40.76$ $42.78$ MMLU $54.12$ $78.76$ $73.55$ $78.17$ MMLU@1[4] $48.01$ $79.48$ $75.23$ $77.90$ MMLU-Pro $24.16$ $53.13$ $61.78$ $67.38$ MMLU-Pro@1[4] $27.13$ $55.76$ $73.21$ $66.96$ GPQA $25.25$ $39.90$ $41.41$ $48.00$ GPQA@1[4] $22.47$ $48.86$ $52.15$ $49.62$ Math Avg $61.38$ $65.33$ $70.16$ $74.93$ Science Avg $34.51$ $57.26$ $58.91$ $64.52$ Science Avg@1[4] $32.54$ $61.37$ $66.86$ $64.83$ Overall $42.81$ 61.32 $65.31$ 68.09
Table 2: Quantitative benchmarks for Nemotron-Nano-12B-v2, showing the impact of RLP. Shaded columns indicate RLP variants; “Post” indicates SFT + RLVR post-training.
Scaling Model Size and Architecture
We further scale RLP to Nemotron-Nano-12B-v2 (Nano, 2025) ($\mathcal{M}_{\mathrm{base}}$), a hybrid Mamba-Transformer language model of 12B parameter size. In this comparison we take an intermediate checkpoint of Nemotron-Nano-12B-v2 trained till $19.8$ trillion tokens and apply RLP for $250$ million tokens only. $\mathcal{M}_{\mathrm{base}}$ on the other hand is trained for $20$ trillion tokens. In addition, we employ an identical post‑training pipeline (SFT $\rightarrow$ RLVR), mirroring the setup used for qwen3-1.7b-base in Table 1. The results as shown in Table 2 confirms that, regardless of model size and families, RLP not only yields a very large improvement at the base stage (Overall 42.81% to 61.32%, a 43% relative gain) but that these gains persist and continue to compound after strong post training. After SFT + RLVR, the RLP trained model improves from 61.32% to 68.09%, maintaining a clear margin over the compute matched baseline (65.31%). The largest relative gains are in scientific reasoning as the Science Avg rises from 34.51% to 57.26% at the base stage and further to 64.52% after post training, compared to 58.91% for the continuously pretrained baseline. This pattern closely mirrors our findings on Qwen3, and demonstrates that RLP scales effectively both to larger parameter counts and to a different architecture family, while remaining compatible with strong downstream alignment. Furthermore, we validate that RLP scales effectively to larger backbones by applying it to qwen3-14b-base, where it improves the overall average from 60.66% to 65.00% after training on 1B tokens, with particularly strong gains in scientific reasoning. Full results for this scaling experiment are provided in the Appendix 12.
@@
Following the experimental setup in RPT (Dong et al., 2025), we trained $\mathcal{M}_{\mathrm{base}}$ on both RPT and RLP methods for one epoch under tokens and flop matched compute budgets before evaluating on our benchmark suite. In the token matched setting, we trained both models on Omni‑MATH (Gao et al., 2024) using the same number of input tokens.
@@
Flop-Matched
$\mathcal{M}_{\mathrm{RPT}}$ $36.66$ $34.38$ $35.68$
$\mathcal{M}_{\textsc{RLP}}$ $45.95$ $38.76$ $42.86$
Table 3: Token- and flop-matched comparisons of RLP and RPT using a qwen3-1.7b-base model.
As we apply RLP to a single token per document while RPT is applied on multiple tokens per document, the number of target tokens for which reward is applied is substantially larger for RPT.
Conversely, for the flop-matched training, both models are trained on Nemotron-CrossThink (as detailed in Appendix 12) for one epoch on the same data, and we ensure that the number of target tokens for which reward is applied is same in both training runs. As summarized in Table Table 3, under Token-Matched setting, RLP achieves uniformly higher aggregates: Math Avg improves by an absolute +2.12% (+4.5% relative), Science Avg by +1.19%(+3.3% relative), and Overall Avg boosts by +1.66% (+4.0% relative). In addition, under the flop matched setting, the improvement is even more prominent. RLP achieves a 20.12% relative improvement on average over RPT. Methodologically, RPT applies reinforcement only to tokens pre‑selected by an auxiliary assistant via entropy filtering and optimizes a sparse, binary next‑token correctness signal that ignores the CoT content, limiting where the signal can be applied. In contrast, RLP evaluates each sampled CoT by the information gain it provides for the observed next token and updates at all positions without an auxiliary filter which yields consistently better averages under the matched setting above. Crucially, this dense, per‑token information‑gain reward supplies richer credit assignment than RPT’s sparse binary signal and, in our matched experiments, empirically yields better performance.
@@ 5 Ablations
Fig. 2 visualizes the trends across three settings: (a) rollouts, (b) completion length, and (c) KL. Please look into §11 for more detailed numbers and per-task breakdowns. More rollouts help up to $G=16$ (Overall $42.17\%$); $G=4$ and $8$ already reach $41.38\%$ and $41.95\%$, while $G=32$ decreases slightly to $41.75\%$ (Fig. 2a). Increasing completion length gives the largest gains. Specifically, overall rises from $11.50\%$ at $64$ to $42.17\%$ at $2048$, with Math/Science moving from $1.12\%/21.88\%$ to $48.06\%/36.29\%$ (Fig. 2b). Extending to $4096$ yields $42.21\%$ at roughly twice the thought budget, so we default to $2048$. Furthermore, a KL anchor does not help. Specifically, $\beta=10^{-4}$ and $10^{-3}$ give $41.35\%$ and $41.44\%$, compared to $42.17\%$ at $\beta=0$, and it also increases memory and step time (Fig. 2c). We therefore use $G=16$, completion length $2048$, and $\beta=0$ in later experiments.